### v1

@@ 2.4 Temperature anisotropy in a magnetized NS
Figure 6: Temperature profiles at different times comparing the analytic solution (solid) and the numerical evolution (stars) of a thermal pulse in a medium embedded in a homogeneous magnetic field. The left (right) panel shows four different times during the evolution of polar (equatorial) profiles in arbitrary units. The simulation has been done with a fully implicit scheme and the linear system is solved with the Thomas algorithm. Figure courtesy of Pérez-Azorín et al (2006).
Figure 7: Temperature anisotropy induced in the NS crust by the presence of a strong magnetic field confined into the crust. The projections of the poloidal field lines are shown with solid lines in the left and right panels, and dashed lines in the central panel. The left panel corresponds to a model without toroidal field, the central panel to a force-free configuration (toroidal magnetic flux contours and poloidal magnetic field lines are aligned), and the right panel shows a model with a toroidal component confined to a narrow region of the crust represented by dashed lines. Figure courtesy of Pérez-Azorín et al (2006).
@@
$T(t,r,\theta)=T_{0}\left(\frac{t_{0}}{t}\right)^{3/2}\exp\left[-\frac{r^{2}}{4t{\kappa^{\perp}}}\left({\sin^{2}\theta}+\frac{\cos^{2}\theta}{1+(\omega_{B}^{e}\tau_{e})^{2}}\right)\right]~,$ (8)
where $T_{0}$ is the central temperature at the initial time $t_{0}$. In Fig. 6 we show the comparison between the analytical (solid) and numerical (stars) solution for a model with $t_{0}=10^{-4}$, $T_{0}=1$, $\kappa^{\perp}=10^{2}$ and $\omega_{B}^{e}\tau_{e}=3$. The boundary conditions employed are $F=0$ at the center and the temperature corresponding to the analytical solution at the surface ($r=1$). Pérez-Azorín et al (2006) found deviations from the analytical solution to be less than 0.1% in any particular cell within the entire domain, even with a relatively low grid resolution of 100 radial zones and 40 angular zones. The same numerical test has been used to test the new 3D code MATINS (Ascenzi et al, 2024).
To conclude this section, the induced anisotropy in a realistic NS reported by Pérez-Azorín et al (2006) is shown in Fig. 7. The figure shows equilibrium thermal solutions, in the absence of heat sources and sinks. The core temperature is kept at $5\times 10^{7}$ K, and the surface boundary condition is given by the $T_{s}-T_{b}$ relation, assuming blackbody radiation. The poloidal component is the same in all models ($B_{p}=10^{13}$ G). The effect of the magnetic field on the temperature distribution can be easily understood by examining the expression of the heat flux (7). When $\omega_{B}^{e}\tau_{e}\gg 1$, the dominant contribution to the flux is parallel to the magnetic field and proportional to $\vec{b}\cdot\vec{\nabla}(e^{\nu}T)$. Thus, in the stationary regime (i.e., $\vec{\nabla}\cdot(e^{2\nu}\vec{F})=0$ if no sources are present), the temperature distribution must be such that $\vec{b}\perp\vec{\nabla}(e^{\nu}T)$: magnetic field lines are tangent to surfaces of constant temperature. This is explicitly visible in the left panel, which corresponds to the stationary solution for a purely poloidal configuration with a core temperature of $5\times 10^{7}$ K. Only near the surface, the large temperature gradient can result in a significant heat flux across the magnetic field lines. When we add a strong toroidal component, the Hall term (the one proportional to $\omega_{B}^{e}\tau_{e}$ in Eq. (7)) results in meridional heat fluxes which lead to a nearly isothermal crust. The central panel shows the temperature distribution for a force-free magnetic field with a global toroidal component, present in both the crust and the envelope. The right panel shows a third model with a strong toroidal component confined to a thin crustal region (dashed lines). It acts as an insulator maintaining a temperature gradient between both sides of the toroidal field.
@@ 3 Magnetic field evolution in the interior of neutron stars: theory review
$\frac{\partial\vec{B}}{\partial t}=-c\vec{\nabla}\times(\mathrm{e}^{\nu}\vec{E})~,$ (9)
@@
In a complete multi-fluid description of plasmas, the set of hydrodynamic equations complements Faraday’s law. From the multi-fluid hydrodynamics equations, a generalized Ohm’s law – in which the electrical conductivity is a tensor – can be derived (Yakovlev and Shalybkov, 1990; Shalybkov and Urpin, 1995)
@@
In some regimes, one can make simplifications to make the problem more affordable (Urpin and Yakovlev, 1980; Jones, 1988; Goldreich and Reisenegger, 1992), although one should incorporate as much physics as possible. The three main processes are Ohmic dissipation, Hall drift (mostly relevant in the crust) and ambipolar diffusion (mostly relevant in the core, Goldreich and Reisenegger 1992; Shalybkov and Urpin 1995; Cumming et al 2004), although additional terms could in principle be included in the induction equation. For instance, there are theoretical arguments proposing additional slow-motion dynamical terms, such as plastic flow (Beloborodov and Levin, 2014; Lander, 2016; Lander and Gourgouliatos, 2019), magnetically induced superfluid flows (Ofengeim and Gusakov, 2018) or vortex buoyancy (Muslimov and Tsygan, 1985; Konenkov and Geppert, 2000; Elfritz et al, 2016; Dommes and Gusakov, 2017). Typically, all these effects are introduced as advective terms, of the type $\vec{E}=-\vec{\mathrm{v}}\times\vec{B}$, with $\vec{v}$ being some effective velocity. The thermoelectric effect (with a contribution to the electric field of the form $\vec{E}=-s\vec{\nabla}T$) has also been proposed to become significant in regions with large temperature gradients (Geppert and Wiebicke, 1991; Wiebicke and Geppert, 1991, 1992, 1995; Geppert and Wiebicke, 1995; Wiebicke and Geppert, 1996), and has been recently revisited in (Gourgouliatos et al, 2022; Gakis and Gourgouliatos, 2024). These additional terms are typically not included in most of the existing literature. However, some of them may play a more important role than expected and should be carefully reconsidered.
@@ 3.1 Ohmic dissipation
$\frac{\partial\vec{B}}{\partial t}+\vec{\nabla}\times\left(\eta\vec{\nabla}\times\vec{B}\right)=0~,$ (12)
@@
In the relaxation time approximation, the electrical conductivity parallel to the magnetic field, $\sigma=e^{2}n_{e}\tau_{e}/m^{*}_{e}$, with $n_{e}$ being the electron number density. Typical values of the electrical conductivity in the crust are $\sigma\sim 10^{22}$$10^{25}$ s-1, several orders of magnitude larger than in the most conductive terrestrial metals described by the band theory in solid state physics. In the core, the even larger electrical conductivity ($\sigma\sim 10^{26}$$10^{29}$ s-1) results in much longer Ohmic timescales, thus potentially affecting the magnetic field evolution only at a very late stage ($t\gtrsim 10^{8}$ yr), when isolated NSs are too cold to be observed. In Fig. 8 we show typical profiles of the electrical conductivity, for the same combinations of $T$ and $B$ shown for the thermal conductivity in Fig. 3. Since, neglecting inelastic scattering, both thermal and electrical conductivities are proportional to the collision time $\tau_{e}$, they share some trends: the suppression of the conduction in the direction orthogonal to a strong magnetic field, and the quantizing effects visible as oscillations around the classical value (Potekhin et al, 2015b; Potekhin and Chabrier, 2018). We note that, if inelastic scattering contributes significantly, $\tau_{e}$ can be different for thermal and electrical conductivities.
@@ 3.2 The Hall drift
Most previous studies of magnetic field evolution in NS crusts (Hollerbach and Rüdiger, 2002, 2004; Pons and Geppert, 2007; Reisenegger et al, 2007; Pons et al, 2009a; Kondić et al, 2011; Viganò et al, 2012, 2013; Gourgouliatos et al, 2013; Marchant et al, 2014; Gourgouliatos and Cumming, 2014b, 2015; Gourgouliatos et al, 2015; Wood and Hollerbach, 2015) have been restricted to 2D simulations. The few recent 3D models (Viganò et al, 2019; Gourgouliatos and Pons, 2019; De Grandis et al, 2021, 2022; Igoshev and Hollerbach, 2023; Igoshev et al, 2023; Dehman et al, 2023c, b) suggest that most 2D features persist, most notably, the role of the Hall term in driving a direct cascade, transferring magnetic energy from large to small scales and thereby enhancing Ohmic dissipation (see Sect. 3.1). Fully 3D simulations, however, reveal distinct behaviors, such as the emergence of long-lived, Hall-driven small-scale azimuthal magnetic structures (see Sect. 6 for details) and the occurrence of an inverse cascade (Brandenburg, 2020; Dehman and Brandenburg, 2025). The latter transfers magnetic energy from small-scale turbulence to larger scales, a process of key relevance for the amplification of large-scale fields in astrophysical contexts.
@@
where $k$ is the wavenumber, and $\chi_{M}(k)$ and $E_{\rm M}(k)$ are the spectra of magnetic helicity and magnetic energy, respectively. Since $\chi_{M}(k)$ may take positive or negative values, the inequality involves its absolute value. Saturation of Eq. (18) at a given $k$ corresponds to maximal helicity at that scale. If this holds for all $k$, the system is said to be maximally helical (Frisch et al, 1975). [^4] A single-mode helical field is also force-free, with $(\vec{\nabla}\times\vec{B})\parallel\vec{B}$ and vanishing Lorentz force. However, a superposition of helical modes with different wavenumbers generally breaks the force-free condition.
For a fully helical field (e.g., $\chi_{M}(k)=2E_{\rm M}(k)/k$ for positive helicity), Frisch et al (1975) showed that neither magnetic energy nor helicity can cascade directly to smaller scales. Instead, nonlinear interactions of modes with wavenumbers $\vec{p}$ and $\vec{q}$ generate new modes with $\vec{k}=\vec{p}+\vec{q}$, constrained such that $|\vec{k}|\leq\max(|\vec{p}|,|\vec{q}|)$ (Brandenburg and Subramanian, 2005). This restriction drives magnetic helicity and energy toward progressively larger length scales in an approximately self-similar fashion, giving rise to the inverse cascade. The role of magnetic helicity in enabling the inverse cascade was first recognized by Frisch et al (1975) and later explored in the context of NSs, initially with box simulations of the Hall term (Brandenburg, 2020) and more recently with global NS models (Dehman and Brandenburg, 2025).
@@ 3.3 The chiral magnetic effect
Beyond macroscopic physics such as Ohmic dissipation, the Hall drift, or dynamos and turbulence during NS formation, some quantum effects can also play a significant role in the magnetic evolution, linking magnetic helicity to fermionic chirality—the handedness of particles (left vs right). One particularly interesting effect is the so-called chiral anomaly, which enables bidirectional transfer between chiral imbalance and magnetic helicity, facilitating the transfer of energy to larger-scale structures in a manner reminiscent of the inverse cascade driven by the non-linear terms in the induction equation (Boyarsky et al, 2012). In this subsection, we briefly review the topic and some recent relevant results.
A small imbalance in the chemical potentials between left- and right-handed electrons, denoted by $\mu_{5}\equiv\mu_{R}-\mu_{L}$ [^5] Note that definitions of $\mu_{5}$ vary in the literature, sometimes differing by a sign Sigl and Leite (2016) or a factor of 2 Kaplan et al (2017), depending on the source., generates an electric current parallel to the magnetic field, an effect known as the Adler-Bell-Jackiw anomaly Adler (1969); Bell and Jackiw (1969). When $\mu_{5}\neq 0$, Maxwell’s equations acquire an additional current term Vilenkin (1980):
@@
The chiral term, which plays a similar mathematical role to turbulent dynamos, has led to studies of stellar core collapse or the proto-NS phase where large-scale fields are generated via chiral asymmetries (Masada et al, 2018; Matsumoto et al, 2022). A major obstacle in these scenarios is the existence of spin-flip processes, driven by the finite electron mass, which quickly suppress the chiral imbalance, reducing the efficiency of the Chiral Magnetic Effect (CME) and inhibiting the Chiral Magnetic Instability (CMI) (Grabowska et al, 2015; Sigl and Leite, 2016; Kamada et al, 2023). This raises questions about the CME’s relevance in such short-lived environments. A recent study by Dehman and Pons (2025) found that in NSs, chiral asymmetries can persist for centuries after birth in the presence of small tangled magnetic structures, enabling a continuous transfer of energy from small to large scales, despite strong suppression by spin-flip processes.
The induction equation must be coupled to the evolution equation for the chiral number density $n_{5}\equiv n_{R}-n_{L}$ (Adler, 1969; Kamada et al, 2023; Dehman and Pons, 2025), which includes both source and sink terms:
@@
Here, the reaction rate $\Gamma_{f}$ accounts for spin-flip interactions, arising from electromagnetic interactions and the finite mass of electrons, and acts as a sink term (Grabowska et al, 2015; Kaplan et al, 2017), while $\Gamma_{w}^{\mathrm{eff}}$ represents the effective weak reaction rate (Epstein and Pethick, 1981) and acts as a source term. The $\vec{E}\cdot\vec{B}$ term governs the coupling between the chiral density and the electromagnetic field: twisting or untwisting magnetic field lines alters the net chirality in the system, acting as either a source or a sink depending on its sign.
Equation (21) should be viewed alongside the time evolution of the magnetic helicity, which can be written in the form Biskamp (1997); Boyarsky et al (2012):
@@
$k_{5}\left(\boldsymbol{x},t\right)=\frac{\left(\vec{\nabla}\times\vec{B}\right)\cdot\vec{B}}{\left(\dfrac{2\mu_{e}^{2}}{m_{e}^{2}c^{4}}\right)\dfrac{B^{2}_{\mathrm{QED}}}{3\pi}+B^{2}}~,$ (24)
@@
The chiral asymmetry can be maintained as long as $\vec{B}\cdot(\vec{\nabla}\times\vec{B})\neq 0$ (non-vanishing current helicity); see Eq. (24). This allows a small but sustained asymmetry to persist, despite the action of spin-flip processes. In this regime, one can insert Eq. (24) into Equation (20), explicitly factoring out the chiral anomaly from the induction equation. In simpler terms, the CME facilitates energy transfer across scales by introducing a new nonlinear term in the magnetic field evolution equation, linked to the current component parallel to the magnetic field $\vec{B}$. We note that this non-linearity is mathematically different from the Hall term, that appears as a quadratic ($B^{2}$) term, driven by the current component perpendicular to the field. In Dehman and Pons (2025) the saturation of the field amplification due to the CME is estimated to be:
@@
In the first detailed simulations of this process, crustal failures were treated in the most simplified manner as star-quakes. By evaluating the accumulated stress, Pons and Perna (2011); Perna and Pons (2011) simulated the frequency and energetics of the internal magnetic rearrangements, which were proposed to be at the origin of magnetar outbursts. This model mimics earthquakes since, under terrestrial conditions, the low densities of the material allow for the propagation of sudden fractures. The Earth mantle, in this respect, can be thought of as brittle. However, materials subject to very slow shearing forces could behave differently and enter a slow plastic flow regime instead. Although the dynamics of failure may differ, the energetic considerations underlying the release of energy due to accumulated magnetic stresses remain broadly similar. Failure or yielding is expected to occur when components of the elastic strain tensor, $\sigma_{ij}$, approach the critical strain threshold, $\sigma_{\mathrm{max}}$, typically in the range of 0.001 to 0.1 Horowitz and Kadau (2009). Assuming that the system evolves slowly and remains close to quasi-equilibrium, the elastic strain can be related to the magnetic stress, ${\cal M}_{ij}$, and the shear modulus, $\mu$, via the relation $\mu\sigma_{ij}=\Delta{\cal M}_{ij}$ (but see discussion at the end of the subsection), where $\Delta$ denotes the change in magnetic stress between the current configuration and the previous equilibrium configuration, established at the time of the previous failure. Making use of this simple algebraic relation, Lander et al (2015) apply the von Mises criterion to establish that the crust would yield when

### v2

@@ 2.4 Temperature anisotropy in a magnetized NS
Figure 6: Temperature profiles at different times comparing the analytic solution (solid) and the numerical evolution (stars) of a thermal pulse in a medium embedded in a homogeneous magnetic field. The left (right) panel shows four different times during the evolution of polar (equatorial) profiles in arbitrary units. The simulation has been done with a fully implicit scheme and the linear system is solved with the Thomas algorithm. Image reproduced with permission from Pérez-Azorín et al. (2006), copyright by ESO.
Figure 7: Temperature anisotropy induced in the NS crust by the presence of a strong magnetic field confined into the crust. The projections of the poloidal field lines are shown with solid lines in the left and right panels, and dashed lines in the central panel. The left panel corresponds to a model without toroidal field, the central panel to a force-free configuration (toroidal magnetic flux contours and poloidal magnetic field lines are aligned), and the right panel shows a model with a toroidal component confined to a narrow region of the crust represented by dashed lines. Image reproduced with permission from Pérez-Azorín et al. (2006), copyright by ESO.
@@
$T(t,r,\theta)=T_{0}\left(\frac{t_{0}}{t}\right)^{3/2}\exp\left[-\frac{r^{2}}{4t{\kappa^{\perp}}}\left({\sin^{2}\theta}+\frac{\cos^{2}\theta}{1+(\omega_{B}^{e}\tau_{e})^{2}}\right)\right]\,$, (8)
where $T_{0}$ is the central temperature at the initial time $t_{0}$. In Fig. 6 we show the comparison between the analytical (solid) and numerical (stars) solution for a model with $t_{0}=10^{-4}$, $T_{0}=1$, $\kappa^{\perp}=10^{2}$ and $\omega_{B}^{e}\tau_{e}=3$. The boundary conditions employed are $F=0$ at the center and the temperature corresponding to the analytical solution at the surface ($r=1$). Pérez-Azorín et al. (2006) found deviations from the analytical solution to be less than 0.1% in any particular cell within the entire domain, even with a relatively low grid resolution of 100 radial zones and 40 angular zones. The same numerical test has been used to test the new 3D code MATINS (Ascenzi et al., 2024).
To conclude this section, the induced anisotropy in a realistic NS reported by Pérez-Azorín et al. (2006) is shown in Fig. 7. The figure shows equilibrium thermal solutions, in the absence of heat sources and sinks. The core temperature is kept at $5\times 10^{7}$ K, and the surface boundary condition is given by the $T_{s}-T_{b}$ relation, assuming blackbody radiation. The poloidal component is the same in all models ($B_{p}=10^{13}$ G). The effect of the magnetic field on the temperature distribution can be easily understood by examining the expression of the heat flux (7). When $\omega_{B}^{e}\tau_{e}\gg 1$, the dominant contribution to the flux is parallel to the magnetic field and proportional to $\vec{b}\cdot\vec{\nabla}(e^{\nu}T)$. Thus, in the stationary regime (i.e., $\vec{\nabla}\cdot(e^{2\nu}\vec{F})=0$ if no sources are present), the temperature distribution must be such that $\vec{b}\perp\vec{\nabla}(e^{\nu}T)$: magnetic field lines are tangent to surfaces of constant temperature. This is explicitly visible in the left panel, which corresponds to the stationary solution for a purely poloidal configuration with a core temperature of $5\times 10^{7}$ K. Only near the surface, the large temperature gradient can result in a significant heat flux across the magnetic field lines. When we add a strong toroidal component, the Hall term (the one proportional to $\omega_{B}^{e}\tau_{e}$ in Eq. (7)) results in meridional heat fluxes which lead to a nearly isothermal crust. The central panel shows the temperature distribution for a force-free magnetic field with a global toroidal component, present in both the crust and the envelope. The right panel shows a third model with a strong toroidal component confined to a thin crustal region (dashed lines). It acts as an insulator maintaining a temperature gradient between both sides of the toroidal field.
@@ 3 Magnetic field evolution in the interior of neutron stars: theory review
$\frac{\partial\vec{B}}{\partial t}=-c\vec{\nabla}\times(\mathrm{e}^{\nu}\vec{E})\,$, (9)
@@
In a complete multi-fluid description of plasmas, the set of hydrodynamic equations complements Faraday’s law. From the multi-fluid hydrodynamics equations, a generalized Ohm’s law — in which the electrical conductivity is a tensor — can be derived (Yakovlev and Shalybkov, 1990; Shalybkov and Urpin, 1995)
@@
In some regimes, one can make simplifications to make the problem more affordable (Urpin and Yakovlev, 1980; Jones, 1988; Goldreich and Reisenegger, 1992), although one should incorporate as much physics as possible. The three main processes are Ohmic dissipation, Hall drift (mostly relevant in the crust) and ambipolar diffusion (mostly relevant in the core, Goldreich and Reisenegger 1992; Shalybkov and Urpin 1995; Cumming et al. 2004), although additional terms could in principle be included in the induction equation. For instance, there are theoretical arguments proposing additional slow-motion dynamical terms, such as plastic flow (Beloborodov and Levin, 2014; Lander, 2016; Lander and Gourgouliatos, 2019), magnetically induced superfluid flows (Ofengeim and Gusakov, 2018) or vortex buoyancy (Muslimov and Tsygan, 1985; Konenkov and Geppert, 2000; Elfritz et al., 2016; Dommes and Gusakov, 2017). Typically, all these effects are introduced as advective terms, of the type $\vec{E}=-\vec{\mathrm{v}}\times\vec{B}$, with $\vec{v}$ being some effective velocity. The thermoelectric effect (with a contribution to the electric field of the form $\vec{E}=-s\vec{\nabla}T$) has also been proposed to become significant in regions with large temperature gradients (Geppert and Wiebicke, 1991; Wiebicke and Geppert, 1991, 1992, 1995; Geppert and Wiebicke, 1995; Wiebicke and Geppert, 1996), and has been recently revisited in Gourgouliatos et al. (2022); Gakis and Gourgouliatos (2024). These additional terms are typically not included in most of the existing literature. However, some of them may play a more important role than expected and should be carefully reconsidered.
@@ 3.1 Ohmic dissipation
$\frac{\partial\vec{B}}{\partial t}+\vec{\nabla}\times\left(\eta\vec{\nabla}\times\vec{B}\right)=0\,$, (12)
@@
In the relaxation time approximation, the electrical conductivity parallel to the magnetic field, $\sigma=e^{2}n_{e}\tau_{e}/m^{*}_{e}$, with $n_{e}$ being the electron number density. Typical values of the electrical conductivity in the crust are $\sigma\sim 10^{22}$$10^{25}$ s-1, several orders of magnitude larger than in the most conductive terrestrial metals described by the band theory in solid state physics. In the core, the even larger electrical conductivity ($\sigma\sim 10^{26}$$10^{29}$ s-1) results in much longer Ohmic timescales, thus potentially affecting the magnetic field evolution only at a very late stage ($t\gtrsim 10^{8}$ yr), when isolated NSs are too cold to be observed. In Fig. 8 we show typical profiles of the electrical conductivity, for the same combinations of $T$ and $B$ shown for the thermal conductivity in Fig. 3. Since, neglecting inelastic scattering, both thermal and electrical conductivities are proportional to the collision time $\tau_{e}$, they share some trends: the suppression of the conduction in the direction orthogonal to a strong magnetic field, and the quantizing effects visible as oscillations around the classical value (Potekhin et al., 2015b; Potekhin and Chabrier, 2018). We note that, if inelastic scattering contributes significantly, $\tau_{e}$ can be different for thermal and electrical conductivities.
@@ 3.2 The Hall drift
Most previous studies of magnetic field evolution in NS crusts (Hollerbach and Rüdiger, 2002, 2004; Pons and Geppert, 2007; Reisenegger et al., 2007; Pons et al., 2009; Kondić et al., 2011; Viganò et al., 2012, 2013; Gourgouliatos et al., 2013; Marchant et al., 2014; Gourgouliatos and Cumming, 2014b, 2015; Gourgouliatos et al., 2015; Wood and Hollerbach, 2015) have been restricted to 2D simulations. The few recent 3D models (Viganò et al., 2019; Gourgouliatos and Pons, 2019; De Grandis et al., 2021, 2022; Igoshev and Hollerbach, 2023; Igoshev et al., 2023; Dehman et al., 2023c, b) suggest that most 2D features persist, most notably, the role of the Hall term in driving a direct cascade, transferring magnetic energy from large to small scales and thereby enhancing Ohmic dissipation (see Sect. 3.1). Fully 3D simulations, however, reveal distinct behaviors, such as the emergence of long-lived, Hall-driven small-scale azimuthal magnetic structures (see Sect. 6 for details) and the occurrence of an inverse cascade (Brandenburg, 2020; Dehman and Brandenburg, 2025). The latter transfers magnetic energy from small-scale turbulence to larger scales, a process of key relevance for the amplification of large-scale fields in astrophysical contexts.
@@
where $k$ is the wavenumber, and $\chi_{M}(k)$ and $E_{\rm M}(k)$ are the spectra of magnetic helicity and magnetic energy, respectively. Since $\chi_{M}(k)$ may take positive or negative values, the inequality involves its absolute value. Saturation of Eq. (18) at a given $k$ corresponds to maximal helicity at that scale. If this holds for all $k$, the system is said to be maximally helical (Frisch et al., 1975). [^4] A single-mode helical field is also force-free, with $(\vec{\nabla}\times\vec{B})\parallel\vec{B}$ and vanishing Lorentz force. However, a superposition of helical modes with different wavenumbers generally breaks the force-free condition.
For a fully helical field (e.g., $\chi_{M}(k)=2E_{\rm M}(k)/k$ for positive helicity), Frisch et al. (1975) showed that neither magnetic energy nor helicity can cascade directly to smaller scales. Instead, nonlinear interactions of modes with wavenumbers $\vec{p}$ and $\vec{q}$ generate new modes with $\vec{k}=\vec{p}+\vec{q}$, constrained such that $|\vec{k}|\leq\max(|\vec{p}|,|\vec{q}|)$ (Brandenburg and Subramanian, 2005). This restriction drives magnetic helicity and energy toward progressively larger length scales in an approximately self-similar fashion, giving rise to the inverse cascade. The role of magnetic helicity in enabling the inverse cascade was first recognized by Frisch et al. (1975) and later explored in the context of NSs, initially with box simulations of the Hall term (Brandenburg, 2020) and more recently with global NS models (Dehman and Brandenburg, 2025).
@@ 3.3 The chiral magnetic effect
Beyond macroscopic physics such as Ohmic dissipation, the Hall drift, or dynamos and turbulence during NS formation, some quantum effects can also play a significant role in the magnetic evolution, linking magnetic helicity to fermionic chirality—the handedness of particles (left vs right). One particularly interesting effect is the so-called chiral anomaly, which enables bidirectional transfer between chiral imbalance and magnetic helicity, facilitating the transfer of energy to larger-scale structures in a manner reminiscent of the inverse cascade driven by the non-linear terms in the induction equation (Boyarsky et al., 2012). In this subsection, we briefly review the topic and some recent relevant results.
A small imbalance in the chemical potentials between left- and right-handed electrons, denoted by $\mu_{5}\equiv\mu_{R}-\mu_{L}$ [^5] Note that definitions of $\mu_{5}$ vary in the literature, sometimes differing by a sign (Sigl and Leite, 2016) or a factor of 2 (Kaplan et al., 2017), depending on the source., generates an electric current parallel to the magnetic field, an effect known as the Adler–Bell–Jackiw anomaly (Adler, 1969; Bell and Jackiw, 1969). When $\mu_{5}\neq 0$, Maxwell’s equations acquire an additional current term (Vilenkin, 1980):
@@
The chiral term, which plays a similar mathematical role to turbulent dynamos, has led to studies of stellar core collapse or the proto-NS phase where large-scale fields are generated via chiral asymmetries (Masada et al., 2018; Matsumoto et al., 2022). A major obstacle in these scenarios is the existence of spin-flip processes, driven by the finite electron mass, which quickly suppress the chiral imbalance, reducing the efficiency of the Chiral Magnetic Effect (CME) and inhibiting the Chiral Magnetic Instability (CMI) (Grabowska et al., 2015; Sigl and Leite, 2016; Kamada et al., 2023). This raises questions about the CME’s relevance in such short-lived environments. A recent study by Dehman and Pons (2025) found that in NSs, chiral asymmetries can persist for centuries after birth in the presence of small tangled magnetic structures, enabling a continuous transfer of energy from small to large scales, despite strong suppression by spin-flip processes.
The induction equation must be coupled to the evolution equation for the chiral number density $n_{5}\equiv n_{R}-n_{L}$ (Adler, 1969; Kamada et al., 2023; Dehman and Pons, 2025), which includes both source and sink terms:
@@
Here, the reaction rate $\Gamma_{f}$ accounts for spin-flip interactions, arising from electromagnetic interactions and the finite mass of electrons, and acts as a sink term (Grabowska et al., 2015; Kaplan et al., 2017), while $\Gamma_{w}^{\mathrm{eff}}$ represents the effective weak reaction rate (Epstein and Pethick, 1981) and acts as a source term. The $\vec{E}\cdot\vec{B}$ term governs the coupling between the chiral density and the electromagnetic field: twisting or untwisting magnetic field lines alters the net chirality in the system, acting as either a source or a sink depending on its sign.
Eq. (21) should be viewed alongside the time evolution of the magnetic helicity, which can be written in the form (Biskamp, 1997; Boyarsky et al., 2012):
@@
$k_{5}\left(\boldsymbol{x},t\right)=\frac{\left(\vec{\nabla}\times\vec{B}\right)\cdot\vec{B}}{\left(\dfrac{2\mu_{e}^{2}}{m_{e}^{2}c^{4}}\right)\dfrac{B^{2}_{\mathrm{QED}}}{3\pi}+B^{2}}\,$, (24)
@@
The chiral asymmetry can be maintained as long as $\vec{B}\cdot(\vec{\nabla}\times\vec{B})\neq 0$ (non-vanishing current helicity); see Eq. (24). This allows a small but sustained asymmetry to persist, despite the action of spin-flip processes. In this regime, one can insert Eq. (24) into Eq. (20), explicitly factoring out the chiral anomaly from the induction equation. In simpler terms, the CME facilitates energy transfer across scales by introducing a new nonlinear term in the magnetic field evolution equation, linked to the current component parallel to the magnetic field $\vec{B}$. We note that this non-linearity is mathematically different from the Hall term, that appears as a quadratic ($B^{2}$) term, driven by the current component perpendicular to the field. In Dehman and Pons (2025) the saturation of the field amplification due to the CME is estimated to be:
@@
In the first detailed simulations of this process, crustal failures were treated in the most simplified manner as star-quakes. By evaluating the accumulated stress, Pons and Perna (2011); Perna and Pons (2011) simulated the frequency and energetics of the internal magnetic rearrangements, which were proposed to be at the origin of magnetar outbursts. This model mimics earthquakes since, under terrestrial conditions, the low densities of the material allow for the propagation of sudden fractures. The Earth mantle, in this respect, can be thought of as brittle. However, materials subject to very slow shearing forces could behave differently and enter a slow plastic flow regime instead. Although the dynamics of failure may differ, the energetic considerations underlying the release of energy due to accumulated magnetic stresses remain broadly similar. Failure or yielding is expected to occur when components of the elastic strain tensor, $\sigma_{ij}$, approach the critical strain threshold, $\sigma_{\mathrm{max}}$, typically in the range of 0.001 to 0.1 (Horowitz and Kadau, 2009). Assuming that the system evolves slowly and remains close to quasi-equilibrium, the elastic strain can be related to the magnetic stress, ${\cal M}_{ij}$, and the shear modulus, $\mu$, via the relation $\mu\sigma_{ij}=\Delta{\cal M}_{ij}$ (but see discussion at the end of the subsection), where $\Delta$ denotes the change in magnetic stress between the current configuration and the previous equilibrium configuration, established at the time of the previous failure. Making use of this simple algebraic relation, Lander et al. (2015) apply the von Mises criterion to establish that the crust would yield when